# Generalized invariants meet constitutive neural networks: A novel framework for hyperelastic materials

### v2

@@ Appendix B Discovered model parameters
Table 6: Model specification for rubber. Discovered material parameters for SI-CANN and GI-CANN in the general strain energy function in (13). The corresponding values of the exponents $\alpha$ and goodness of git are given in Table 3.
@@
Table 7: Model specification for brain cortex. Discovered material parameters for SI-CANN and GI-CANN in the general strain energy function (13). The corresponding values of the exponents $\alpha$ and goodness of git are given in Table 4.
@@
$w^{\mathrm{id}}_{1}$[MPa] $\mathcal{J}_{\alpha_{+}}-3]$ 0.997 - - - - -
$w^{\mathrm{id}}_{2}$ [MPa] $\mathcal{J}_{\alpha_{-}}-3] - - -$ 0.009 - -
$w^{\mathrm{exp}}_{1,1}[-],\;w^{\mathrm{exp}}_{1,2}$ [MPa] $\exp([\mathcal{J}_{\alpha_{+}}-3])-1$ - 0.703, 0.626 - - - -
$w^{\mathrm{exp}}_{2,1}[-],\;w^{\mathrm{exp}}_{2,2}$ [MPa] $\exp([\mathcal{J}_{\alpha_{-}}-3])-1$ - - - - - 0.096, 0.096
$w^{\mathrm{ln}}_{2,1}[-],\;w^{\mathrm{ln}}_{2,2}$ [MPa] $\ln(1-[\mathcal{J}_{\alpha_{-}}-3])$ - - 1.090, 0.921 - 1.064, 0.941 -

### v3

@@ Appendix B Discovered model parameters
Table 6: Model specification for rubber. Discovered material parameters for SI-CANN and GI-CANN in the general strain energy function in (13). The corresponding values of the exponents $\alpha$ and goodness of fit are given in Table 3.
@@
Table 7: Model specification for brain cortex. Discovered material parameters for SI-CANN and GI-CANN in the general strain energy function (13). The corresponding values of the exponents $\alpha$ and goodness of fit are given in Table 4.
@@
$w^{\mathrm{id}}_{1}$[kPa] $\mathcal{J}_{\alpha_{+}}-3]$ 0.997 - - - - -
$w^{\mathrm{id}}_{2}$ [kPa] $\mathcal{J}_{\alpha_{-}}-3] - - -$ 0.009 - -
$w^{\mathrm{exp}}_{1,1}[-],\;w^{\mathrm{exp}}_{1,2}$ [kPa] $\exp([\mathcal{J}_{\alpha_{+}}-3])-1$ - 0.703, 0.626 - - - -
$w^{\mathrm{exp}}_{2,1}[-],\;w^{\mathrm{exp}}_{2,2}$ [kPa] $\exp([\mathcal{J}_{\alpha_{-}}-3])-1$ - - - - - 0.096, 0.096
$w^{\mathrm{ln}}_{1,1}[-],\;w^{\mathrm{ln}}_{1,2}$ [kPa] $\ln(1-[\mathcal{J}_{\alpha_{+}}-3])$ - - - - - -
$w^{\mathrm{ln}}_{2,1}[-],\;w^{\mathrm{ln}}_{2,2}$ [kPa] $\ln(1-[\mathcal{J}_{\alpha_{-}}-3])$ - - 1.090, 0.921 - 1.064, 0.941 -